# Multi-Agent Trajectory Prediction with Difficulty-Guided Feature Enhancement Network

### v2

@@ III-E Difficulty-guided decoder
To capture features of plausible future trajectories in the scene, we introduce a Difficulty-Guided Decoder to obtain the future trajectories of agents that are relatively easier to predict. Firstly, we decode the actor features $\hat{\mathcal{X}}^{(l)}$ in scence-centric, which have undergone multiple layers of feature interactions. Each agent receives six predicted trajectories. Following the ADAPT [4] trajectory decoder to get higher quality decoded trajectories by predicting endpoints for refinement. The decoding process can be expressed as follows:
^$\hat{\mathcal{E}}=\mathbf{MLP}_{\text{end}}(\hat{\mathcal{X}}^{(l)})+\mathbf{MLP}_{\text{refine}}(\text{cat}[\hat{\mathcal{X}}^{(l)},\hat{\mathcal{E}}])$ (15)
$\hat{\mathcal{P}}=\text{cat}[\mathbf{MLP}_{\text{traj}}(\text{cat}[\hat{\mathcal{X}}^{(l)},\hat{\mathcal{E}}]),\hat{\mathcal{E}}]$ (16)
@@
where endpoints $\hat{\mathcal{E}}$ are predicted and refined by concatenating with the agent features $\hat{\mathcal{X}}^{(l)}$, then the trajectories $\hat{\mathcal{P}}$ are predicted and merged with the refined endpoints.
@@
7:        $\hat{\mathcal{P}}_{masked}$.add($\hat{\mathcal{P}}_{l}$)
@@
The Argoverse 1 dataset show that the predicted trajectories of most of the vehicles in the scenarios exhibit a high degree of concentration, with more than 90% of the samples of straight ahead trajectories in each case [21]. This suggests that the motion patterns of most agents can be easily captured, reflecting real-world traffic scenarios. In order to prevent higher difficulty prediction trajectories from entering the subsequent modules and generating cumulative errors, we introduced a difficulty masker to filter the initial prediction results. ScenceTransformer[22], inspired by recent approaches to language modeling, has pioneered the idea of using a masking strategy as a query to the model, enabling it to invoke a single model to predict agent behavior in multiple ways (marginal or joint prediction). This mechanism in this paper is intended to mask out the initial prediction trajectories of the more difficult agents, thus providing better control over the prediction accuracy. This process can be formalized as Algo. 1.
III-F Future feature enhancement
For the reliable future trajectories we first flatten them by performing the flatten operation on multiple future trajectories for feature dimension alignment. Then future trajectory feature extraction is performed by a simple MLP layer. These two operations can be formulated as:
^$\hat{\mathcal{F}}=\mathbf{MLP}(\mathbf{flatten}(\hat{\mathcal{P}}_{\text{masked}}))$ (17)
Subsequently, we interact the future trajectory feature $\hat{F}$ with the agent-centric interactive actor feature $\hat{\mathcal{A}}^{(k)}$ via multi-head attention. This operation aims to impose certain constraints on the prediction of other agents through the reliable future trajectory features. Finally, we need to perform one last interaction using the original map features $\hat{\mathcal{M}}$. This step is crucial as it aims to refocus on the reachable lane lines of the map after imposing social constraints through the future trajectory features[23]. These two operations can be formulated as:
$\mathcal{H}=\mathbf{MHA}_{AF}(\hat{\mathcal{A}}^{(k)},\hat{\mathcal{F}})$ (18)
$\hat{\mathcal{O}}=\mathbf{MHA}_{AM}(\mathcal{H},\hat{M})$ (19)
III-G Final prediction
We perform feature fusion operations on the final features $\hat{O}$ and the scence-centric feature $\hat{X}^{(l)}$, expanding the decoding dimension to 256. Our final predictor can generate multimodal trajectories for all agents in the scene and their corresponding probabilities in a single inference. This process can be described as follows:
${\mathcal{A}}_{fusion}=\text{cat}(\hat{\mathcal{O}},\hat{\mathcal{X}}^{(l)})$ (20)
$\mathcal{E}=\mathbf{MLP}_{\text{end}}({\mathcal{A}}_{fusion})+\mathbf{MLP}_{\text{refine}}(\text{cat}[{\mathcal{A}}_{fusion},\mathcal{E}])$ (21)
$\mathcal{K}=\mathbf{softmax}(\mathbf{MLP}_{\text{cls}}(\text{cat}[\mathcal{A}_{fusion},\mathcal{E}]))$ (22)
$\mathcal{P}=\text{cat}[\mathbf{MLP}_{\text{traj}}(\text{cat}[{\mathcal{A}}_{fusion},\mathcal{E}]),\mathcal{E}]$ (23)
here, $\mathcal{P}$ represents the multimodal prediction trajectory and its corresponding probability $\mathcal{K}$ that we finally obtain.
Table I: Comparisons with the state-of-the-art methods listed on the leaderboard of the Argoverse 1 motion Forecasting test set. The numbers highlighted in bold represent the best-performing results among them.
@@ III-G Final prediction
Model Training. Although our feature enhancement component outputs trajectories during the intermediate process, our training process remains end-to-end. Our DGFNet obtains initial and final predictions in a single pass. Similar to the previous method [16, 17, 1, 13], we supervise the output trajectories $P$ and $\hat{\mathcal{P}}$ during the training process by smoothing the $\mathcal{L}1$ loss, which is expressed as ${\mathcal{L}}_{reg}$ and ${\mathcal{L}}_{reg}^{c}$. For the categorization loss in $\mathcal{K}$, we use the maximum marginal loss, which is expressed as ${\mathcal{L}}_{cls}$. Our overall loss expression is as follows:
@@
IV Experiments
IV-A Experimental setup
@@
Datasets. Our model DGFNet is tested and evaluated on a very challenging and widely used self-driving motion prediction dataset: the Argoverse 1&2 [32, 33]. Both motion prediction datasets provide agent tracking trajectories and semantically rich map information at a frequency of 10Hz over a specified time interval. The prediction task in Argoverse 1 is to predict trajectories for the next 3 seconds based on the previous 2 seconds of historical data. The dataset contains a total of 324,557 vehicle trajectories of interest extracted from over 1000 hours of driving. Argoverse 2 contains 250,000 of the most challenging scenarios officially filtered from the self-driving test fleet. Argoverse 2 predicts the next 6 seconds from the first 5 seconds of historical trajectory data. To ensure fair comparisons between models, both datasets were subjected to official data partitioning and the test set was evaluated using the Eval online test server.
@@
Evaluation Metrics. We have adopted the standard testing and evaluation methodology used in motion prediction competitions to assess prediction performance. Key metrics for individual agents include Probabilistic minimum Final Displacement Error (p-minFDE), Minimum Final Displacement Error (minFDE), Minimum Average Displacement Error (minADE), Miss Rate (MR) and Drivable Area Compliance (DAC). Where p-minFDE, MR, and minFDE reflect the accuracy of the predicted endpoints, and minADE indicates the overall bias in the predicted trajectories. DAC, reflects the compliance of the predicted outcomes.
@@
Implementation Details. We generate lane vectors for lanes that are more than 50 meters away from any available agent. The number of layers $l$ and $k$ of interaction features are set to 3. All layers except the final decoder have 128 output feature channels. In addition, the weight parameters in the loss function are set to $\alpha$=0.7, $\beta$=0.1 and $\lambda$=0.2. The hyperparameter $\tau$ in the difficulty masker was set to 5. The model was trained 50 epochs on an Nvidia RTX 3090 with a batch size of 16. For the experiments on Argoverse 1&2, we used a segmented constant learning rate strategy: up to the 5th epoch we used $5\times 10^{-5}$; from the 6th epoch to the 40th, we used $5\times 10^{-4}$; and thereafter, we used $5\times 10^{-5}$ until training was complete.
IV-B Quantitative Results
@@
Quantitative evaluations on the Argoverse 1&2 datasets demonstrate that DGFNet exhibits significant performance advantages across various metrics compared to state-of-the-art methods. Our method remains competitive even when compared to ensemble strategies used by other top models like QCNet and ProphNet. Through an ablation study, we validate the effectiveness of scene feature fusion, future feature interaction, and a difficulty masking mechanism, resulting in substantial improvements in prediction accuracy. Computational efficiency analysis shows that DGFNet achieves high prediction accuracy with relatively fewer parameters, ensuring its practical applicability in real-time autonomous driving scenarios.
Comparison with State-of-the-Arts in Agoverse 1. We conducted performance comparisons on the Argoverse 1 test set using both single-model and multi-model approaches between our method and the most classical and latest state-of-the-art methods. Generally, ensemble methods involve using K-means to match the closest trajectories from multiple models, followed by weighted computation to obtain the final result. We obtained 5 models by setting different initial random seeds and performed ensemble inference using the aforementioned method. The results are shown in Tab.I. Compared to single-model methods, our approach consistently ranks at the top in almost all metrics. Even when compared with methods using ensemble strategies, our approach remains highly competitive, achieving prediction accuracy comparable to the leading models such as QCNet and ProphNet in the leaderboard. However, our method has advantages in terms of parameter size and inference time, which will be detailed in the following sections.
@@
Table II: Comparisons with the state-of-the-art methods listed on the leaderboard of the Argoverse 2 single agent Forecasting.
@@
Comparison with State-of-the-Arts in Agoverse 2. In our comparison of performance with several recent single-model inference methods on the Argoverse 2 test set, we found that while the advantage of our method is not as pronounced as in Argoverse 1, it remains reasonably acceptable. We speculate that this is primarily due to the increased prediction horizon in the Argoverse 2 dataset, which has led to less robust initial prediction results. Moreover, due to the inclusion of categories such as pedestrians and cyclists in Argoverse 2, the effectiveness of the parameter settings for our difficulty masker has also been impacted. Overall, our model demonstrates highly competitive performance within an acceptable range of model parameters across both datasets.
Table III: Comparison of prediction performance (minFDE) on high-difficulty samples in the Argoverse 1 validation set.
@@
Comparison of Prediction Performance on High-difficulty Samples. Tab.III shows the highest minFDE errors of each method on the Argoverse 1 validation set. We reasoned on the validation set with LaneGCN’s pre-trained model to obtain some sample sets with the highest errors, and then reasoned on these sample sets using our base model and DGFNet, respectively. It can be observed that the average accuracy difference for all samples is minimal compared to Our base. However, our method, DGFNet, has significantly lower minFDE errors across all percentage ranges, clearly demonstrating the remarkable advantage of our method in mitigating the long-tail problem in trajectory prediction.
@@
Table IV: The ablation study results of different components on the Argoverse 1 validation set.
DM FFI SFF p-minFDE $\downarrow$ minADE $\downarrow$ minFDE $\downarrow$
@@
$\times$ $\times$ $\times$ 1.559 0.657 0.941
@@
1.492 0.632 0.892
Table V: Performance versus $\tau$ parameter of DGFNet evaluated on the Argoverse 1 validation set.
@@
IV-C Analysis and Discussion
Ablation Study. We conducted an ablation study on the Argoverse 1 validation set. The initial model only used agent-centric historical trajectories and maps for feature extraction. The second model incorporated scene-centric features into the initial model, achieving some performance improvement through Scene Feature Fusion (SFF). The third model added the Future Feature Interaction (FFI) module, enabling interaction between extracted historical features and future features. The final model used the Difficulty Masker (DM) to mask out future trajectories that are difficult to predict, ensuring that only reliable trajectories are used for future interaction. Of concern is that without adding Difficulty Masker, the performance of direct future enhancement drops dramatically, failing because the second stage model may find shortcuts during optimization if the results given in the first stage are good enough [34]. According to Tab.IV, comparing the base model with DGFNet, the improvement in predictive performance is evident. And we also did ablation experiments on the effect of parameter $\tau$ on the model. According to Tab.V, the model performance obtained by setting parameter $\tau$ to 5 is the best.
Table VI: Comparison on Computational performance with other methods.
Method p-minFDE $\downarrow$ Param(M) $\downarrow$ Infer time(ms) $\downarrow$
(K=6)
LaneGCN[16] 2.054 3.68 17.2
GANet[3] 1.789 5.21 25.3
QCNet* [31] 1.693 7.30 -
HiVT-128[13] 1.842 2.47 13.2
LAformer[25] 1.835 2.59 42.5
FFINet*[6] 1.729 6.20 -
Prophnet*[5] 1.694 15.20 -
DGFNet 1.693 4.53 32.1

### v3

@@ II-E Trajectory Decoder
Difficulty-Guided decoder. To capture features of plausible future trajectories in the scene, we introduce a Difficulty-Guided Decoder to obtain the future trajectories of agents that are relatively easier to predict. Firstly, we decode the actor features $\hat{\mathcal{X}}^{(l)}$ in scene-centric representation, which have undergone multiple layers of feature interactions. Each agent receives six predicted trajectories. Following the ADAPT [4] trajectory decoder to get higher quality decoded trajectories by predicting endpoints for refinement. The decoding process is described as follows:
^$\begin{split}&\hat{\mathcal{E}}=\rm{MLP}_{end}(\hat{\mathcal{X}}^{(l)})+\rm{MLP}_{refine}([\hat{\mathcal{X}}^{(l)},\hat{\mathcal{E}}]),\\ &\hat{\mathcal{P}}=[\rm MLP_{traj}([\hat{\mathcal{X}}^{(l)},\hat{\mathcal{E}}]),\hat{\mathcal{E}}]\end{split}$ (8)
@@
where endpoints $\hat{\mathcal{E}}$ are predicted and refined by concatenating with the actor features $\hat{\mathcal{X}}^{(l)}$, the trajectories $\hat{\mathcal{P}}$ predicted with the refined endpoints and []delimited-[][$\cdot$][ ⋅ ] represents the concatenation.
@@
7:        $\hat{\mathcal{P}}_{masked}$.add($\hat{\mathcal{P}}_{i}$)
@@
Table I: Comparisons with the methods listed on the leaderboard of the Argoverse 1 motion Forecasting test set. The numbers highlighted in bold represent the best-performing results among them.
@@
The Argoverse 1 dataset shows that the predicted trajectories of most of the vehicles in the scenarios exhibit a high degree of concentration, with more than 90% of the samples of straight ahead trajectories in each case [33]. This suggests that the motion patterns of most agents can be easily captured, reflecting real-world traffic scenarios. In order to prevent higher difficulty prediction trajectories from entering the subsequent modules and generating cumulative errors, we introduced a difficulty masker to filter the initial prediction results. SceneTransformer[34], inspired by recent approaches to language modeling, has pioneered the idea of using a masking strategy as a query to the model, enabling it to invoke a single model to predict agent behavior in multiple ways (marginal or joint prediction). This mechanism in this paper is intended to mask out the initial prediction trajectories of the more difficult agents, thus providing better control over the prediction accuracy. This process can be formalized as Algo. 1.
Future feature enhancement. For the reliable future trajectories, we first flatten them by performing the flatten operation on multiple future trajectories for feature dimension alignment. Then future trajectory feature extraction is performed by a simple MLP layer. These two operations can be described as follows:
^$\hat{\mathcal{F}}=\rm{MLP}(\rm{Flatten}(\hat{\mathcal{P}}_{masked}))$ (9)
Subsequently, we interact the future trajectory features $\hat{F}$ with the agent-centric interactive actor features $\hat{\mathcal{A}}^{(k)}$ via multi-head attention. This operation aims to impose certain constraints on the prediction of other agents through the reliable future trajectory features. Finally, we need to perform one last interaction using the original map features $\hat{\mathcal{M}}$. This step is crucial as it aims to refocus on the reachable lane lines of the map after imposing social constraints through future trajectory features[35]. These two operations can be described as follows:
$\begin{split}&\mathcal{H}=\rm{MHA}_{AF}(\hat{\mathcal{A}}^{(k)},\hat{\mathcal{F}}),\\ &\hat{\mathcal{O}}=\rm{MHA}_{AM}(\mathcal{H},\hat{\mathcal{M}})\end{split}$ (10)
Final decoder. We perform feature fusion operations on the final features $\hat{O}$ and the scene-centric feature $\hat{X}^{(l)}$, expanding the decoding dimension to 256. Our final predictor can generate multimodal trajectories for all agents in the scene and their corresponding probabilities in a single inference. This process is described as follows:
$\begin{split}&{\mathcal{A}}_{fusion}=[\hat{\mathcal{O}},\hat{\mathcal{X}}^{(l)}],\\ &\mathcal{E}=\rm{MLP}_{end}({\mathcal{A}}_{fusion})+\rm{MLP}_{refine}([{\mathcal{A}}_{fusion},\mathcal{E}]),\\ &\mathcal{K}=\rm{Softmax}(\rm{MLP}_{cls}([\mathcal{A}_{fusion},\mathcal{E}])),\\ &\mathcal{P}=[\rm{MLP}_{traj}([{\mathcal{A}}_{fusion},\mathcal{E}]),\mathcal{E}]\end{split}$ (11)
where $\mathcal{P}$ represents the multimodal prediction trajectory, its corresponding probability $\mathcal{K}$ that we finally obtain and []delimited-[][$\cdot$][ ⋅ ] represents the concatenation.
Model training. Although our feature enhancement component outputs trajectories during the intermediate process, our training process remains end-to-end. Our model obtains initial and final predictions in a single pass. Similar to the previous method [21, 22, 1, 18], we supervise the output trajectories $\mathcal{P}$ and $\hat{\mathcal{P}}$ during the training process by smoothing the $\mathcal{L}1$ loss, which is expressed as ${\mathcal{L}}_{reg}$ and ${\mathcal{L}}_{reg}^{c}$. For the categorization loss in $\mathcal{K}$, we use the maximum marginal loss, which is expressed as ${\mathcal{L}}_{cls}$. Our overall loss expression is as follows:
@@
III Experiments
III-A Experimental Setup
@@
Datasets. Our model is tested and evaluated on a very challenging and widely used self-driving motion prediction dataset: the Argoverse 1&2 [36, 37]. Both motion prediction datasets provide agent tracking trajectories and semantically rich map information at a frequency of 10Hz over a specified time interval. The prediction task in Argoverse 1 is to predict trajectories for the next 3 seconds based on the previous 2 seconds of historical data. The dataset contains a total of 324,557 vehicle trajectories of interest extracted from over 1000 hours of driving. Argoverse 2 contains 250,000 of the most challenging scenarios officially filtered from the self-driving test fleet. Argoverse 2 predicts the next 6 seconds from the first 5 seconds of historical trajectory data. To ensure fair comparisons between models, both datasets were subjected to official data partitioning and the test set was evaluated using the Eval online test server.
@@
Evaluation metrics. We have adopted the standard testing and evaluation methodology used in motion prediction competitions to assess prediction performance. Key metrics for individual agents include Probabilistic minimum Final Displacement Error (p-minFDE), Minimum Final Displacement Error (minFDE), Minimum Average Displacement Error (minADE), Miss Rate (MR) and Drivable Area Compliance (DAC). Where p-minFDE, MR, and minFDE reflect the accuracy of the predicted endpoints, and minADE indicates the overall bias in the predicted trajectories. DAC reflects the compliance of the predicted outcomes. The key metrics for multiple agents are obtained by averaging the performance metrics of the labeled agents in each scenario of the test dataset. It is worth noting that these labeled agents share the same scene prediction probability.
@@
Implementation details. We generate lane vectors for lanes that are more than 50 meters away from any available agent. The number of layers $l$ and $k$ of interaction features are both set to 3. All layers except the Final Decoder have 128 output feature channels. In addition, the weight parameters in the loss function are set to $\alpha$=0.7, $\beta$=0.1 and $\lambda$=0.2. The hyperparameter $\tau$ in the difficulty masker was set to 5. The model was trained 50 epochs on an Nvidia RTX 3090 with a batch size of 16. For the experiments on Argoverse 1&2, we used a segmented constant learning rate strategy: up to the 5th epoch we used $5\times 10^{-5}$; from the 6th epoch to the 40th, we used $5\times 10^{-4}$; and thereafter, we used $5\times 10^{-5}$ until training was complete.
III-B Quantitative Results
@@
Quantitative evaluations on the Argoverse 1&2 datasets demonstrate that DGFNet exhibits significant performance advantages across various metrics compared to SOTA methods. Our method remains competitive even when compared to ensemble strategies used by other top models like QCNet[32] and ProphNet[5]. Through ablation experiments, we demonstrate the effectiveness of each module mechanism and the experiments on adjusting the relevant parameters. Computational efficiency analysis shows that DGFNet achieves high prediction accuracy with relatively fewer parameters, ensuring its practical applicability in real-time autonomous driving scenarios.
@@
Table II: Comparisons with the methods listed on the leaderboard of the Argoverse 2 single-agent Forecasting.
@@
Table III: Comparisons with the methods listed on the leaderboard of the Argoverse 2 muti-world Forecasting.
Method Year avgP-MinFDE $\downarrow$ avgMinADE $\downarrow$ avgMinFDE $\downarrow$ actorMR $\downarrow$
(K=6) (K=6) (K=6) (K=6)
LaneGCN[21] 2020 3.90 1.49 3.24 0.37
HiVT[18] 2022 2.85 0.88 2.20 0.26
FJMP[38] 2023 2.59 0.81 1.89 0.23
FFINet[6] 2024 2.44 0.77 1.77 0.24
DGFNet - 2.37 0.73 1.68 0.20
Comparison with prediction performance in Agoverse 1. We conducted performance comparisons on the Argoverse 1 test set using both single-model and multi-model approaches between our method and the most classical and latest methods. Generally, ensemble methods involve using K-means to match the closest trajectories from multiple models, followed by weighted computation to obtain the final result. We obtained 5 models by setting different initial random seeds and performed ensemble inference using the aforementioned method. The results are shown in Tab.I. Compared to single-model methods, our approach consistently ranks at the top in almost all metrics. Even when compared with methods using ensemble strategies, our approach remains highly competitive, achieving prediction accuracy comparable to the leading models such as QCNet[32] and ProphNet[5] in the leaderboard. However, our method has advantages in terms of parameter size and inference time, which will be detailed in the following sections.
Comparison with prediction performance in Agoverse 2. As shown in Tab.II & Tab.III , in our comparison of several trajectory prediction methods on the Argoverse 2 single-agent and multi-agent test sets, we found that while the advantage of our method is not as pronounced as in Argoverse 1, it remains reasonably acceptable. We speculate that this is primarily due to the increased prediction horizon in the Argoverse 2 dataset, which has led to less robust initial prediction results. Moreover, due to the inclusion of categories such as pedestrians and cyclists in Argoverse 2, the effectiveness of the parameter settings for our difficulty masker has also been impacted. Overall, our model demonstrates highly competitive performance within an acceptable range of model parameters across both datasets.
Table IV: Comparison of prediction performance (minFDE $\downarrow$) on high-difficulty samples in the Argoverse 1 validation set.
@@
Comparison of prediction performance on high-difficulty samples. Tab.IV shows the highest minFDE errors of each method on the Argoverse 1 validation set. We reasoned on the validation set with LaneGCN’s[21] pre-trained model to obtain some sample sets with the highest errors and then reasoned on these sample sets using our Agent-centric base model and DGFNet. It can be observed that the average accuracy difference for all samples is minimal compared to our base model. However, our method has significantly lower minFDE errors across all percentage ranges, clearly demonstrating the remarkable advantage of our method in mitigating the long-tail problem in trajectory prediction.
@@
Table V: The ablation study results of different components on the Argoverse 1 validation set.
Base Model DM FFI SFF p-minFDE $\downarrow$ minADE $\downarrow$ minFDE $\downarrow$
@@
Scene-Base $\times$ $\times$ $\times$ 1.641 0.699 1.044
Agent-Base $\times$ $\times$ $\times$ 1.559 0.657 0.941
@@
- $\times$ 1.550 0.652 0.934
DGFNet 1.492 0.632 0.892
Table VI: Performance versus loss fuction weight parameter of DGFNet evaluated on the Argoverse 1 validation set.
$\alpha$ $\lambda$ p-minFDE $\downarrow$ minADE $\downarrow$ minFDE $\downarrow$
(K=6) (K=6) (K=6)
0.5 0.4 1.519 0.640 0.911
0.6 0.3 1.502 0.635 0.900
0.7 0.2 1.492 0.632 0.892
0.8 0.1 1.503 0.636 0.902
Table VII: Performance versus $\tau$ parameter of DGFNet evaluated on the Argoverse 1 validation set.
@@
III-C Analysis and Discussion
Ablation study. We conducted an ablation study on the Argoverse 1 validation set. As shown in V, our base model only utilizes agent-centric or scene-centric historical trajectories and maps for feature extraction and interaction. The third model performs Scene Feature Fusion (SFF) on agent-centric and scene-centric features, achieving a certain level of performance improvement. The fourth model incorporates the Future Feature Interaction (FFI) module, enabling interaction and alignment between the extracted historical features, original map features, and future features. The fifth model removes the SFF module and introduces the Difficulty Masker (DM), where the filtered future trajectory features are directly embedded without undergoing the interaction process.
$\begin{split}&\rm{SceneBase}=\rm{FiDecoder}(\hat{\mathcal{X}}^{(l)}),\\ &\rm{AgentBase}=\rm{FiDecoder}(\hat{\mathcal{A}}^{(k)}),\\ &\rm{w/\>SFF}=\rm{FiDecoder}([\hat{\mathcal{A}}^{(k)},\hat{\mathcal{X}}^{(l)}]),\\ &\rm{w/\>SFF\&DM}=\rm{FiDecoder}([\hat{\mathcal{A}}^{(k)},\hat{\mathcal{F}}]),\\ \end{split}$ (13)
where $\rm{SceneBase}$ represents the prediction results obtained by our scene-centric base model, $\rm{AgentBase}$ represents the prediction results obtained by our agnet-centric base model, $\rm{FiDecoder}$ represents the entire Final Decoder module, $\rm{w/\>SFF}$ and $\rm{w/\>SFF\&DM}$ represent the prediction results of the base model with $\rm{SFF}$ and $\rm{SFF\&DM}$ added, respectively. []delimited-[][$\cdot$][ ⋅ ] represents the concatenation.
Of concern is that the performance of the fifth model drops significantly, because the second stage model may find shortcuts during optimization if the results given in the first stage are good enough [39].
Regarding the selection of loss function weights, based on our previous experience in adjusting parameters on the base model and the principle that the most important weight should take the largest proportion, we first fixed the weight of ${\mathcal{L}}_{cls}$ ($\beta$) to 0.1. Then, we conducted ablation experiments by gradually increasing the weight of ${\mathcal{L}}_{reg}$ ($\alpha$) starting from 0.5. As shown in VI, we found that the model performance was quite good when $\alpha=0.7$, $\beta=0.1$, and $\lambda=0.2$.
We also did ablation experiments on the effect of parameter $\tau$ on the model. According to Tab.VII, the model performance obtained by setting parameter $\tau$ to 5 is the best.
Figure 4: The x-axis represents the predicted result p-minFDE, and the y-axis represents the model’s parameter size. Both values should be minimized for better performance.